OpenStudy (anonymous):

the top three languages in nevada are english, spanish, and tagalog. the population of nevada is approx. 2,117,000. if 3/5 of the people speak spanish or tagalog or both, how many people speak english?

OpenStudy (anonymous):

assuming no one speaks both, it would be \[\frac{2}{5}\times 2,117,000\]

OpenStudy (anonymous):

so no one speaks english in nevada (what I wrote)

OpenStudy (mathmate):

Assuming those who speak spanish or tagalog or both do NOT speak English, then English speakers = (2/5)*population.

OpenStudy (anonymous):

thanks :D this is great help! :)

OpenStudy (mathmate):

You're welcome! Hope it helped. :) In fact, I find the question not very clear. It is not clear to me if the question asks for English as a native language (mother tongue) or just English speaker? Many people speak multiple languages.

OpenStudy (anonymous):

that's so true where are you from? @mathmate

OpenStudy (mathmate):

Canada

OpenStudy (anonymous):

United states....

OpenStudy (anonymous):

canada's second language is french right?

OpenStudy (mathmate):

In fact, French is one of Canada's two official languages (beside English), however, French speakers make up only about 25% of Canada's population.

OpenStudy (anonymous):

wow cool!

OpenStudy (mathmate):

Yep, but it's cold up here! :D

OpenStudy (anonymous):

wow in chicago it's not even snowing

OpenStudy (mathmate):

You're close to the lake, so it's mild in winter, and cool in summer.

OpenStudy (anonymous):

yup :D